What frequency of sound traveling in air at 20 degrees C has a wavelength equal to 1.6 m, the average height of a person? (Hz)


1
Expert's answer
2021-10-06T08:17:33-0400

Let's first find the speed of sound in air at "20^{\circ}C":


"v=(331.3+0.606\cdot\theta)=(331.3+0.606\cdot20^{\circ}C)=343.42\ \dfrac{m}{s}."

Finally, we can find the frequency of the sound from the wave-speed formula:


"v=f\lambda,""f=\dfrac{v}{\lambda}=\dfrac{343.42\ \dfrac{m}{s}}{1.6\ m}=215\ Hz."